You bought a call option of K=490 and it expires in 1 month. Current stock price is 489. You paid

$5. If the stock price on the expiration day is 500. Do you make money or lose money, how

much do you make or lose? What if the stock price is 480, or less than 480?
Answer #1

A call option makes money when spot price at expiry is more than strike price

1.
Profit=MAX(500-490,0)-5=5
Profit of $5

2.
Profit=MAX(480-490,0)-5=-5
Loss of $5

3.
Loss of $5